Treatment-Free Beekeeping by Clive & Shân Hudson 2020 as published in BBKA News No. 227 July 2020. Includes a concluding remark by Professor Thomas Seeley.
An earlier version of this article was published in Welsh Beekeeper 206, Winter 2019.
Our bee monitor was recently featured in Raspberry Pi’s magazine and on their blog: ‘BeeMonitor’, The MagPi 94 (June 2020).
Available on the Raspberry Pi Blog: https://www.raspberrypi.org/blog/monitoring-bees-with-a-raspberry-pi-and-beemonitor/
Available as a pdf: https://magpi.raspberrypi.org/issues/94/pdf
This article is an update of our experience of Treatment Free beekeeping.
11 Years of Treatment Free Beekeeping by Clive & Shan Hudson as published in Welsh Beekeeper 206, Winter 2019.
“Varroa Has Lost its Sting – Our Experience of Treatment Free Beekeeping”
Presentation given at Welsh Beekeepers’ Convention at the Royal Welsh Showground, Builth Wells on 30th March 2019 to an audience of approximately 165. This presentation is similar to that given at the BIBBA Confernce 2018, but differs in two main aspects:
- given from the ‘Our Experience’ rather than the ‘My Experience’ viewpoint to emphasise the Treatment Free (TF) approach being taken by many beekeepers as well as ourselves.
- includes updated data and map of TF Beekeepers in North-West Wales.
Following the presentation, I was joined by Joe and Chris Ibbertson for a lively Q&A session.
Varroa Has Lost its Sting – Our Experience of Treatment Free Beekeeping – WBKA Conference 2019, Builth Wells
This article is by Claudia Blauert, from Germany following her visit to Wales, and views our Treatment Free beekeeping from a German perspective.
“Varroa Has Lost its Sting – My Experience of Treatment Free Beekeeping”
Presentation given at BIBBA Conference 6th – 9th September 2018, Royal Agricultural University, Cirencester. Following the presentation, and after being joined by Joe and Chris Ibbertson, there was time for a Q&A. The audience of around 60 indicated by a show of hands that approximately half were already beekeeping Treatment Free.
Should we continue treating our bees? Or is it time to put our trust in nature?
- Varroa Has Lost its Sting by Clive & Shan Hudson 2016 as published in BBKA News No. 223 December 2016 (also Welsh BeeKeeper Summer /Autumn 2016 No. 193 & 194)
- Wild Honey Bees Of The Glaslyn Estuary, Snowdonia by Clive & Shan Hudson, Autumn 2015 as published in The Welsh Beekeeper 190, Autumn 2015.
- Notes on Treatment Free Beekeeping by Clive & Shan Hudson. Supporting evidence for Treatment Free Beekeeping and the evolution of honey bees to co-exist with Varroa is accumulating locally, nationally and internationally. These are our January 2018 notes on this topic.
Event Log (April 2012 – January 2013):
- 21st April 2012: hive opened for 6 minutes at 14.30 hrs. with an ambient temperature of 12°C. Frames were removed to check the sensor in the Cluster.The Cluster temperature soon dropped to 27°C, but recovered to its tolerance temperature range within minutes of the hive being reassembled.
- 3rd May 2012: Honey super was added above the brood box, position for sensor T3 was changed from ‘Above Crown’ to a new position in the middle of the honey super.
- 17th May 2012: hive was opened for inspection at 12.30 hrs. The ambient temperature was 16°C and the hive was open for about 8 minutes. Bees were loosely clustered in the centre of the Honey Super adjacent to the temperature sensor,(see photo); the wire seen below the wire queen excluder goes to the Cluster sensor.The honey super was temporarily removed to allow access to the brood frames. Half the brood frames were inspected for queen cells. The brood looked good, the queen was seen; there were no queen cells. The graph shows a comparatively small dip in the temperature recorded for the Honey Super and the Cluster.This small drop in temperature was corrected in minutes of the hive being reassembled.
- 23rd May 2012: hive opened at 15:45hrs. Ambient temperature was about 18°C and the hive was open for 12 minutes; between crown-board coming off and going back on. Nine out of the eleven frames had brood which looked good and the queen was seen. The honey super had lots of bees and freshly stored honey, a second super was added. The Cluster temperature and the Honey Super dipped a little during the inspection, but soon recovered. The relatively hot weather this week has produced a lot of bee activity. Some of our hives have swarmed, as have others in the area. We believe the high temperature in the Honey Super, mainly over 30°C, shows a ‘honey flow’ for the last three days since May 21st. How long will it last?
- 26th May 2012: The high temperatures and honey flow continue. A third honey super was added today. The second super added three days ago is now full of bees.The Honey Super sensor remains in its original position in the first super, which is now full of honey!
- 12th June 2012: Hive opened for full inspection. Bees in all three supers. The lowest super with the Honey Super sensor was heavy with honey! (See photo).The brood box had a number of queen cups, some with eggs and at least one queen cell with royal jelly and a young larvae. The Cluster sensor was in a patch of open cells with eggs; beyond the eggs was an arc of sealed brood (see photo). The plan is to allow the colony to swarm and to see if this can be related to any temperature changes. The steep temperature rise in the Honey Super is probably related to the heavy super being ‘half dropped’ and receiving a hefty jolt on the hive stand; the bees were not happy about this and clearly got ‘hot and bothered’! Is this new knowledge?
- 25th June 2012 12.00 hrs: Hive opened for full inspection, the brood box is mainly full of sealed brood, some (older) eggs and the queen were seen. There were a number of queen cells with torn side openings and a few new queen cells with larvae. The hive opening is shown with a dip in the ‘Cluster’ temperature, which the bees soon return to ‘normal’ after the hive is reassembled.
- 25th June 2012 15.15 hrs: The colony swarmed! This corresponded to a peak temperature in the cluster of over 36°C. This is the first time the cluster temperature has gone above 36°C since monitoring started on 25th March. The swarm settled in a damson tree 10m away from the hive and after 20min returned to the hive. (As this is an experimental hive, a box with a queen excluder had been fitted over the hive entrance after 12th June.)
- 30th June 2012: Poor weather prevented intervention prior to this date. The weather was warmer in the afternoon today and the graph suggests the hive may have tried to swarm again; we were not there to see the hive.
- 1st July 2012: We opened the hive today and split the colony as a swarm prevention measure. We left one frame of the emerging brood, the queen, flying bees and the super of honey on the monitored hive. The remaining space in the brood box was filled with frames of foundation. The queen excluder box was removed from the enterance.
- 4th July 2012: The hive swarmed again at 16:00 hrs! Luckily the swarm settled not far from the hive and near to the ground. The marked queen was spotted and put in a plastic cage. The cage was placed adjacent to the hive enterance where the swarm reformed. The hive was opened and the frame of brood removed. The ‘Cluster’ sensor was refixed on a different frame that had empty pulled comb (see photo). The queen was released into the brood box, the swarm gradually returned into the hive.
- 5th July 2012 :A feeder with sugar syrup was put on above the super at 20.45hrs. This seemed to relate to an increased temperature rise in the Honey Super and Cluster, however, the Ambient temperature also rose during the evening.
- 15th September 2012: 17.00hrs. The two supers on the hive had a mix of frames with sealed honey and empty frames; these were rearranged. The super with empty frames was placed above the queen excluder on the brood box and the Honey Super sensor was repositioned onto a central empty frame. An ‘escape board’ was placed on top and the super with honey frames put above. Clearly related to the above manipulation, the graph shows a dip in the temperature of both the Cluster and Honey Super, relating to the hive being opened and the two supers temporarily removed.The Cluster temperature soon recovers. Following on what is particularly interesting at this event is the pronounced rise in the Honey Super temperature of 16C (from 17C-33C) over the following 2-3 hours. The temperature then stayed high for some 6 hours, before gradually returning to 20C. My guess for this reaction is ‘stress’ caused by the disruption to the bees, ie the moving of the supers and frames, and the imposition of the clearer board. The hypothesis is disruption equals stress, equals rapid temperature increase?? (See Event Log for 12th June for a comparable event.)
- 27th September 2012: Miller type feeder placed directly on the hive brood box with 4kg of sugar syrup. As the super had few bees these were brushed out and the super was placed above the feeder, i.e. the super was no longer in contact with the bees, and being ‘stored’ with the sensor still in place.
- 15th October 2012: Hive opened for two reasons: Firstly, to remove the feeder, secondly, out of curiosity, because for the last five weeks or so the Cluster temperature has gradually been dropping from its summer norm of about 34C to nearer 30C. The queen was seen, along with eggs, larvae and sealed brood. The brood area is now significantly smaller than in the summer and restricted to the lower half of the frames. See photograph. This has left the Cluster sensor near the top edge of the brood area and explains why the Cluster temperature is lower. The super was replaced above a queen excluder, above the brood box.
- 6th January 2013: No, the colony has not started to lay! This is an experimental hive and it was opened today, the first time since 15th October 2012, to check on the state of the colony. The main event is shown on the graph by the steep rise in the temperature of the cluster. The bees were not aggressive but they were – not surprisingly – agitated by the disturbance. This caused the Cluster temperature to rise by 10°C in a very short period of time. This also increased the temperature in the super which is not occupied by the cluster. There was no brood. The area of brood seen on 15th October now contained cells filled with liquid honey over which the bees were clustered. There was still a good arc of sealed honey above the bees.The yellow marked queen was seen; can you see her on the attached (rather poor) photograph?